OpenStudy (anonymous):

Solve for x: ln(3x)^2=16.

OpenStudy (anonymous):

ln(3x)=4 - The range of ln(x) does not include the negative root of 16, -4, so we only use the positive root. e^ln(3x)=e^4, x=(e^4)/3

OpenStudy (anonymous):

Thanks.